OpenStudy (anonymous):

Write a quadratic variation equation if g(x) varies directly with x2, and g(x) = 108 when x = 6.

OpenStudy (jmark):

Given g(x) varies directly with x2, and g(x) = 108 when x = 6. let =>g(x) = Kx^2 =>108 = K6^2 =>K = 108/6^2 = 3 =>g(x) = 3x^2 then the quadratic equation is g(x)=3x^2 CBSE board date sheet http://goo.gl/9swacQ